Find the elasticity of y with respect to x for the following: a.y^(6)=x^(5)quad b.(y)/(x)=(x+1)^(a)(y-1)^(b)quad (a and b are constants)


1
Expert's answer
2021-05-17T09:57:58-0400
"E=y^/_x\frac{x}{y}"

a.


"E=\frac{5x^4}{6y^5}\times \frac{x}{x^5}=\frac{5}{6y^5}"

b.


"y^/=\frac{(x+1)^a(a-x-1)(y-1)}{x^2(y-1)^b(y-1-by)}"

"E=\frac{(a-x-1)(y-1)}{(y-1)^{2b}(y-1-by)}"